a spring toy weighing 1kg on a spring balance suddenly jump upwards a bit standing near the toy notices that the scale of the balance reads 1.05kg in this process mwximum acceleration of toy is

Grade:12

1 Answers

Arun
6 years ago
a=(1.05-1)×10=0.05×10=0.5m/s^2Hence it will be the maximum acceleration of the toy...........RegardsArun (askIITians forum expert)